OpenStudy (anonymous):

Find the slope, If it exists, of the line containing the pair of points. (4,6) and (5,-9) The slope m=?

OpenStudy (anonymous):

m = -15

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) m=-9-6/5-4 m=-15/1 m=-15

OpenStudy (anonymous):

slope formula -9-6 over 5-4 -15 1 -15 divided by 1 is =-15